# The Big-O Problem for Max-Plus Automata is Decidable (PSPACE-Complete)

### v2

@@ 1. Introduction
Chistikov, Kiefer, Murawski and Purser [CKMP20, CKMP22] study the big-O problem over the non-negative rational semiring, where it is shown to be undecidable in general, but decidable for certain restrictions on the ambiguity or the accepted language. Similarly, in [CLM+22], two restrictions of the big-O problem, also known to be undecidable in general, are studied on specific subclasses: the boundedness problem (a.k.a. limitedness), and the zero isolation problem [^1] The boundedness problem is the special case of the big-O problem when $[\![\mathcal{B}]\!]=1$. The zero isolation problem is the special case when $[\![\mathcal{A}]\!]=1$. Chistikov et al. [CKMP22] deal with the problem whether $[\![\mathcal{A}]\!]\leq c[\![\mathcal{B}]\!]$ so without $+c$, but the two problems are equivalent (see Remark 4).
@@
The paper expands on [DP23], giving full proofs and explanations. In Section 2, we give the definition of max-plus automata and introduce a running example that we will use all along the paper. In Section 3, we state the big-O problem and a simplified version of it, prove their PSPACE-hardness and reduce our result to prove that the simplified big-O problem is PSPACE (Theorem 2). In Section 4, we give a high-level description of this proof and define the semigroups that will be used and the stabilisation and flattening operations. In Section 5, we define witnesses and give a decision procedure that we show to be PSPACE. Sections 6, 7, 8 and 9 are then dedicated to prove that the decision procedure is sound and complete. The organisation and content of these sections are explained at the end of Section 5, when the suitable notions have been introduced. Finally, in Section 10, we provide a sequence of max-plus automata that require increasingly intricate witnesses (for a notion defined in the paper); this section is new compared with [DP23].
@@ 3. Decidability of the big-O problem
Proposition 3.1 states that it is enough to prove that the simplified big-O problem is in PSPACE in order to conclude the proof of Theorem 1. Proposition 3.3 gives the hardness parts of Theorems 1 and 2 (since the simplified problem is a particular instance of the general one).
@@ 4.1. Proof schema
Up to now the semigroup of asymptotic behaviour witnesses unboundedness of sequences of words, but not yet the difference in growth rates in $\mathcal{A}$ and $\mathcal{B}$. This is where our proof gives new insight. We introduce a second operation, which we call flattening, and behaves as follows: consider an element $(p,x,p,M)$ in the semigroup of asymptotic behaviour such that $x=\infty$, so witnessing some sequence of words with unbounded weights in $\mathcal{A}$ from $p$ to $p$. We would like to iterate the words in this sequence, no longer looking at unboundedness, but rather at which entries in $M$ will grow linearly as the elements of the sequence are repeated. We define the flattening operation to do exactly that. After flattening, we obtain an element $(p,x,p,M^{\prime})$ witnessing a sequence of words $(w_{i})_{i}$ such that it is possible for an entry in $M^{\prime}$ to have value $1$, corresponds to paths with unbounded weights, but only if the asymptotic growth rate of these weights is little-o of the respective weights in $\mathcal{A}$ from $p$ to $p$. In other words, unbounded runs of maximal growth rate continue to be represented by $\infty$, but runs which are not of the maximal growth rate are ‘flattened’ back to $1$, even if they are unbounded.
@@ 4.2. Semigroup of paths
  • $(p,x,q,M)\otimes(p^{\prime},x^{\prime},q^{\prime},M^{\prime})=(p,x\otimes x^{\prime},q^{\prime},M\otimes M^{\prime})$ if $q=p^{\prime}$ and $\bot$ otherwise.
@@
  • $(p,x,q,M)\otimes(p^{\prime},x^{\prime},q^{\prime},M^{\prime})=(p,x\otimes x^{\prime},q^{\prime},M\otimes M^{\prime})$ if $q=p^{\prime}$ and $\bot$ otherwise.
@@ 6.1. Factorisation trees
A factorisation tree on $w$ is a finite ordered tree in which every node $\nu$ in the tree is labelled by a element in $\overline{\mathfrak{M}}_{\mathcal{A},\mathcal{B}}$, denoted $\alpha(\nu)$, such that:
@@ 6.2. Contributors
Proposition 6.1.
Given a word $w$ such that $[\![\mathcal{B}]\!](w)\neq-\infty$, every node in any factorisation tree on $w$ has a non empty set of contributors.
Proof 6.2.
@@ 7. No tree has a fault implies Big-O
The last inequality holds only if the set of contributors $C_{\nu}$ is not empty, which is the case by Proposition 6.1 since we assume $[\![\mathcal{B}]\!](w)\geq 0$ for all $w\in\Sigma^{*}$.
@@
Let $\nu_{1},\dots,\nu_{d}$ be the children of $\nu$. By definition and idempotency, $\overline{\mathsf{val}_{\mathcal{A}}(\nu)}=\overline{\mathsf{val}_{\mathcal{A}}(\nu_{m})}$ for all $m=1,\ldots,d$. If $\overline{\mathsf{val}_{\mathcal{A}}(\nu)}=0$, then we directly get the result since the set of contributors $C_{\nu}$ is not empty by Proposition 6.1 (because we assume that $[\![\mathcal{B}]\!](w)\geq 0$ for all $w\in\Sigma^{*}$). Let’s suppose now that $\overline{\mathsf{val}_{\mathcal{A}}(\nu)}=\overline{\mathsf{val}_{\mathcal{A}}(\nu_{m})}=1$ for all $m=1,\ldots,d$.
@@ 10.1. Construction of $\mathcal{A}_{n}$ and $\mathcal{B}_{n}$
Let us first show how we can use Lemmas 23 and 24 to conclude the proof. We prove by induction that for all $m\leq k<n$,

### v4

@@ 1. Introduction
Chistikov, Kiefer, Murawski and Purser [CKMP20, CKMP22] study the big-O problem over the non-negative rational semiring, where it is shown to be undecidable in general, but decidable for certain restrictions on the ambiguity or the accepted language. Similarly, in [CLM+22], two restrictions of the big-O problem, also known to be undecidable in general, are studied on specific subclasses: the boundedness problem (a.k.a. limitedness), and the zero isolation problem. [^1] The boundedness problem is the special case of the big-O problem when $[\![\mathcal{B}]\!]=1$. The zero isolation problem is the special case when $[\![\mathcal{A}]\!]=1$. Chistikov et al. [CKMP22] deal with the problem whether $[\![\mathcal{A}]\!]\leq c[\![\mathcal{B}]\!]$ so without $+c$, but the two problems are equivalent (see Remark 6).
@@
The paper expands on [DP23], giving full proofs and explanations. In Section 2, we give the definition of max-plus automata and introduce a running example that we will use all along the paper. In Section 3, we state the big-O problem and a simplified version of it, prove their PSPACE-hardness and reduce our result to prove that the simplified big-O problem is in PSPACE (Theorem 2). In Section 4, we give a high-level description of this proof and define the semigroups that will be used and the stabilisation and flattening operations. In Section 5, we define witnesses and give a decision procedure that we show to be PSPACE. Sections 6, 7, 8 and 9 are then dedicated to prove that the decision procedure is sound and complete. The organisation and content of these sections are explained at the end of Section 5, when the suitable notions have been introduced. Finally, in Section 10, we provide a sequence of max-plus automata that require increasingly intricate witnesses (for a notion defined in the paper); this section is new compared with [DP23].
@@ 3. Decidability of the big-O problem
Proposition 3 states that it is enough to prove that the simplified big-O problem is in PSPACE in order to conclude the proof of Theorem 1. Proposition 4 gives the hardness parts of Theorems 1 and 2 (since the simplified problem is a particular instance of the general one).
@@ 4.1. Proof schema
Up to now the semigroup of asymptotic behaviour witnesses unboundedness of sequences of words, but not yet the difference in growth rates in $\mathcal{A}$ and $\mathcal{B}$. This is where our proof gives new insight. We introduce a second operation, which we call flattening, and behaves as follows: consider an element $(p,x,p,M)$ in the semigroup of asymptotic behaviour such that $x=\infty$, so witnessing some sequence of words with unbounded weights in $\mathcal{A}$ from $p$ to $p$. We would like to iterate the words in this sequence, no longer looking at unboundedness, but rather at which entries in $M$ will grow linearly as the elements of the sequence are repeated. We define the flattening operation to do exactly that. After flattening, we obtain an element $(p,x,p,M^{\prime})$ witnessing a sequence of words $(w_{i})_{i}$ such that it is possible for an entry in $M^{\prime}$ to have value $1$, corresponding to paths with unbounded weights, but only if the asymptotic growth rate of these weights is little-o of the respective weights in $\mathcal{A}$ from $p$ to $p$. In other words, unbounded runs of maximal growth rate continue to be represented by $\infty$, but runs which are not of the maximal growth rate are ‘flattened’ back to $1$, even if they are unbounded.
@@ 4.2. Semigroup of paths
  • $(p,x,q,M)\otimes(p^{\prime},x^{\prime},q^{\prime},M^{\prime})=(p,x+x^{\prime},q^{\prime},M\otimes M^{\prime})$ if $q=p^{\prime}$ and $\bot$ otherwise.
@@
  • $(p,x,q,M)\otimes(p^{\prime},x^{\prime},q^{\prime},M^{\prime})=(p,x+x^{\prime},q^{\prime},M\otimes M^{\prime})$ if $q=p^{\prime}$ and $\bot$ otherwise.
@@ 6.1. Factorisation trees
A factorisation tree on $w$ is a finite ordered tree in which every node $\nu$ in the tree is labelled by an element in $\overline{\mathfrak{M}}_{\mathcal{A},\mathcal{B}}$, denoted $\alpha(\nu)$, such that:
@@ 6.2. Contributors
Proposition 14.
Given a word $w$ such that $[\![\mathcal{B}]\!](w)\neq-\infty$, every node in any factorisation tree on $w$ has a non non-empty of contributors.
Proof 6.1.
@@ 7. No tree has a fault implies Big-O
The last inequality holds only if the set of contributors $C_{\nu}$ is not empty, which is the case by Proposition 14 since we assume $[\![\mathcal{B}]\!](w)\geq 0$ for all $w\in\Sigma^{*}$.
@@
Let $\nu_{1},\dots,\nu_{d}$ be the children of $\nu$. By definition and idempotency, $\overline{\mathsf{val}_{\mathcal{A}}(\nu)}=\overline{\mathsf{val}_{\mathcal{A}}(\nu_{m})}$ for all $m=1,\ldots,d$. If $\overline{\mathsf{val}_{\mathcal{A}}(\nu)}=0$, then we directly get the result since the set of contributors $C_{\nu}$ is not empty by Proposition 14 (because we assume that $[\![\mathcal{B}]\!](w)\geq 0$ for all $w\in\Sigma^{*}$). Let’s suppose now that $\overline{\mathsf{val}_{\mathcal{A}}(\nu)}=\overline{\mathsf{val}_{\mathcal{A}}(\nu_{m})}=1$ for all $m=1,\ldots,d$.
@@ 10.1. Construction of $\mathcal{A}_{n}$ and $\mathcal{B}_{n}$
Let us give a rough intuition of what automaton $\mathcal{B}_{n}$ is computing on words of the shape:
$(a_{n}(a_{n-1}(\ldots(a_{2}(a_{1}^{\ell_{1}})a_{2})^{\ell_{2}}\ldots a_{n-1})^{\ell_{n-1}}a_{n})^{\ell_{n}}$
  • $\mathcal{B}_{1}=\mathcal{B}_{1,1}$ computes the function that maps each word to 00.
  • $\mathcal{B}_{2}=\max(\mathcal{B}_{2,1},\mathcal{B}_{2,2})=\mathcal{B}$ from the running example (up to renaming of letters) and computes the function that maps each word to the maximum between the longest block of $a_{1}$’s and the number of $a_{2}$’s.
  • $\mathcal{B}_{3}=\max(\mathcal{B}_{3,1},\mathcal{B}_{3,2},\mathcal{B}_{3,3})$ computes a function that maps each word of the shape $(a_{3}(a_{2}(a_{1}^{\ell_{1}})a_{2})^{\ell_{2}}a_{3})^{\ell_{3}}$ to the maximum of the three following components:
    • (for $\mathcal{B}_{3,1}$) the number of $a_{2}$’s and $a_{3}$’s, so big-O of $\ell_{2}\ell_{3}$,
    • (for $\mathcal{B}_{3,2}$) between all pairs of consecutive $a_{3}$’s, take the largest number of $a_{1}$’s appearing between two $a_{2}$’s, and then take the sum, so it would give big-O of $\ell_{1}\ell_{3}$,
    • (for $\mathcal{B}_{3,3}$) the largest number of $a_{1}$’s appearing between two $a_{3}$’s (ignoring $a_{2}$’s), so big-O of $\ell_{1}\ell_{2}$.
  • for the general case, $\mathcal{B}_{n}$ computes a function that maps each word of the shape
    $(a_{n}(a_{n-1}(\ldots(a_{2}(a_{1}^{\ell_{1}})a_{2})^{\ell_{2}}\ldots a_{n-1})^{\ell_{n-1}}a_{n})^{\ell_{n}}$
    to the maximum of the $n$ following components:
    • (for $\mathcal{B}_{n,1}$) the number of $a_{2}$’s, $a_{3}$’s,… and $a_{n}$’s, so big-O of $\ell_{2}\ell_{3}\dotsm\ell_{n}$,
    • (for $\mathcal{B}_{n,m}$ with $1<m\leq n$) between all pairs of consecutive letters $a_{i}$’s and $a_{j}$’s with $i,j>m$, take the largest number of $a_{1}$’s appearing between two $a_{m}$’s (ignoring the other letters), and then take the sum, so it would give big-O of $\ell_{1}\ell_{2}\dotsm\ell_{m-1}\ell_{m+1}\dotsm\ell_{n}$.
@@
Let us first show how we can use Lemmas 26 and 27 to conclude the proof. We prove by induction that for all $m\leq k<n$,